OpenStudy (anonymous):

Find a model for simple harmonic motion if the position at t = 0 is 2, the amplitude is 2 centimeters, and the period is 6 seconds.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Do you have any idea how to start this? For example, do you know what is the amplitude in each option?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

i need alot of help

OpenStudy (mertsj):

They have told you that the ordered pair (0,2) is on the graph.

OpenStudy (mertsj):

In the first choice, you have 6sin(2t)

OpenStudy (anonymous):

yes (:

OpenStudy (mertsj):

Replace t with 0 and evaluate. Do you get 2?

OpenStudy (anonymous):

So, a simple harmonic motion (or wave) has the following form.\[d = A \cos(\frac{2π}{T}t+k) = A \cos(ωt+k)\] or \[d = A \sin(\frac{2π}{T}t+k) = A \cos(ωt+k)\] where A is the amplitude, T is the period and ω is the angular frequency. So from the formula above, what do you think is the amplitude for the choice A?

OpenStudy (anonymous):

Forgot to mention that k is the initial phase of the harmoniv motion.

OpenStudy (anonymous):

no i get 0

OpenStudy (mertsj):

Then move on to the next option. Replace t with 0 and see if you get 2.

OpenStudy (anonymous):

for b i got square root of 3 if i plug it in.

OpenStudy (mertsj):

So move on to the next choice.

OpenStudy (anonymous):

for c i got -1

OpenStudy (mertsj):

Actually, in the second choice, you get 0 again because pi/3 times 0 is 0 and the sin of 0 is 0

OpenStudy (mertsj):

Are you remembering that anything times 0 is 0?

OpenStudy (anonymous):

for D i got like 1.9

OpenStudy (anonymous):

and yeah i do remember.

OpenStudy (mertsj):

D: 2 cos(6t)

OpenStudy (mertsj):

What is 6 times 0 ?

OpenStudy (anonymous):

0

OpenStudy (mertsj):

What is the cosine of 0 ?

OpenStudy (anonymous):

1

OpenStudy (mertsj):

What is 2 times 1 ?

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Now explain to me again how you got 1.9 for D ?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2cos%2812%29

OpenStudy (anonymous):